Caldwell Home Appliances Inc. is estimating the activity cost associated with producing ovens and refrigerators. The indirect labor can be traced into four separate activity pools, based on time records provided by the employees. The budgeted activity cost and activity-base information are provided as follows:

Activity

Activity Pool Cost

Activity Base

Procurement

$12,600

Number of purchase orders

Scheduling

90,000

Number of production orders

Materials handling

11,000

Number of moves

Product development

50,000

Number of engineering changes

Total cost

$163,600

The estimated activity-base usage and unit information for two product lines was determined as follows:

Number of

Purchase Orders

Number of

Production Orders

Number of

Moves

Number of

Engineering Change Orders

Units

Ovens

400

800

300

80

1,000

Refrigerators

300

400

200

120

500

Totals

700

1,200

500

200

1,500

a. Determine the activity rate for each activity cost pool.

Procurement (per purchase order)

$

Scheduling (per production order)

$

Materials handling (per move)

$

Product development (per engineering change order)

$

b. Determine the activity-based cost per unit of each product. Round the per unit rates to the nearest cent.

Ovens

$ per unit

Refrigerators

$ per unit

Please show calculations
Answer #1
a.) Activity rate for each activity pool Amount $
Procurement (per purchase order) 18 =12600/700
Scheduling (per production order) 75 =90000/1200
Materials handling (per move) 22 =11000/500
Product development (per engineering change order) 250 =50000/200
b.) Amount $
Ovens Refrigerators
Procurement         7,200                5,400
Scheduling       60,000              30,000
Materials handling         6,600                4,400
Product development       20,000              30,000
Total activity cost allocated      93,800              69,800
Activity based cost per unit of each product
Ovens $ 93.80 =93800/1000
Refrigerators $ 139.60 =69800/500
Working:
Amount $
Ovens Refrigerators
Procurement =18*400 =18*300
Scheduling =75*800 =75*400
Materials handling =22*300 =22*200
Product development =250*80 =250*120
Total activity cost allocated      93,800              69,800